Solve the given equation and identify it as a conditional equation, an inconsistent equation, or an identity.
5(p – 2) – 9(p + 9) = 5 – 4p

Answer by stanbon(75887)

Solve the given equation and identify it as a conditional equation, an inconsistent equation, or an identity.
5(p – 2) – 9(p + 9) = 5 – 4p
5p - 10 -9p -81 = -35
-4p - 91 = -35
4p = -56
p = -14
============
The equation is conditional; it is true when p = -14.
============
Cheers,
Stan H.